OpenStudy (anonymous):

(3x^3)^2 a.9x^6 b. 9x^5 c. 6x^5 d. 27x^6

OpenStudy (campbell_st):

the power of 2 operates on both 3 and x^3 the index rule for power of a power is \[(x^a)^b = x^{a \times b}\] so really you have \[(3x^3)^2 = 3^2 \times (x^3)^2\] hope this helps

OpenStudy (anonymous):

ya it does but do you add or multiply the exponents

OpenStudy (campbell_st):

read the rule above... its called power of a power...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(x-2)^2

OpenStudy (campbell_st):

squaring something means to multiply something by itself so you have (x-2)*(x -2) which can be written as x(x-2) -2(x -2) just disrtibute and collect like terms.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(2+sqrt of 3)(2-sqrtof 3)